生体熱力学 2 単位 (選択)

Biochemical Thermodynamics  松木均 · 教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

【授業目的】生体関連物質が組織化して構築される分子集合体の熱力学的な取り 扱いおよびそれら集合体への環境変数 (温度，圧力，添加物) の影響について 講述する． 【授業概要】本講義の前半部分では，生体関連物質 (界面活性剤，脂質や両親媒性 薬物) が自己会合して形成する分子集合体 (単分子膜，ミセルおよびベシクル) の取り扱いを熱力学的観点から解説する．後半部分では，これらの分子集合 体の示す様々な性質および環境変数 (温度，圧力，添加物) による集合体の構 造変化について講述する．さらに分子集合体の工学的応用や工業的利用につ いても触れる．本科目は，工業に関する科目である． 【授業形式】講義 【キーワード】生体関連物質，分子集合体，熱力学，相挙動，圧力 【先行科目】『物理化学 2』 (1.0)，『生物物理化学 2』 (1.0) 【関連科目】『生物物理化学特論』 (0.5) 【履修要件】学部教育における物理化学および生物物理化学を理解していること． 【履修上の注意】 [注意] 【到達目標】 1. 生体関連物質が形成する分子集合体の熱力学的取り扱いを理解する． 2. 生体関連物質集合体の性質と環境変数の影響について理解する． 【授業計画】 1. 分子集合体の熱力学 (1) 吸着の界面熱力学 1:表面および界面張力 2. 分子集合体の熱力学 (2) 吸着の界面熱力学 2:吸着膜の取り扱い 3. 分子集合体の熱力学 (3) 単分子膜の相転移 1:不溶性単分子膜 4. 分子集合体の熱力学 (4) 単分子膜の相転移 2:可溶性単分子膜:可溶性単分子膜 5. 分子集合体の熱力学 (5) 会合体形成の熱力学 1:相分離モデル 6. 分子集合体の熱力学 (6) 会合体形成の熱力学 2:質量作用モデル 7. 分子集合体の熱力学 (7) 希薄溶液におけるミセル形成 1:臨界ミセル濃度 8. 分子集合体の熱力学 (8) 希薄溶液におけるミセル形成 2:影響因子 9. 分子集合体の構造と機能 (1) 溶解挙動 1:クラフト点 10. 分子集合体の構造と機能 (2) 溶解挙動 2:曇点 11. 分子集合体の構造と機能 (3) 二分子膜の相挙動 1:飽和リン脂質 12. 分子集合体の構造と機能 (4) 二分子膜の相挙動 2:不飽和リン脂質 13. 分子集合体の構造と機能 (5) 二分子膜への圧力効果 1:転移の体積変化 14. 分子集合体の構造と機能 (6) 二分子膜への圧力効果 2:圧力誘起指組み構造

形成

15. 分子集合体の構造と機能 (7) 二分子膜混合系の相挙動:ドメイン形成

16. 総括，質疑，レポート作成

【成績評価基準】出席率 80%以上で，課題レポート (100%) で評価する．【教科書】授業中に適宜資料を配布する. 【参考書】授業中に紹介する．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216731 【対象学生】他学科学生も履修可能【連絡先】

⇒ 松木 (化生棟 607, 088-656-7513, matsuki@bio.tokushima-u.ac.jp) ���� (オ

フィスアワー: 金曜日 16:20-17:50) 【備考】

⋄ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

⋄ 到達目標 1 は授業計画 1-8 に，到達目標 2 は授業計画 9-15 に関係する．

Biochemical Thermodynamics 2 units (selection)

Hitoshi Matsuki · Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering

Target⟩ Thermodynamic treatments for molecular aggregates constructed by selfassociation of bio-related substances and effects of environmental variables such as temperature, pressure and additives on the aggregates are described.

Outline⟩ The former part of this lecture reviews treatments for molecular aggregates such as monolayers, micelles and vesicles constructed by self-association of bio-related substances like surfactants, lipids and amphiphilic drugs from a thermodynamic viewpoint. The latter part describes various nature of these aggregates and structure changes of the aggregates by environmental variables such as temperature, pressure and additives. Further industrial application and utilization of molecular aggregates are also explained. This lecture deals with industrial subjects.

Style⟩ Lecture �������⟩ bio-related substance, molecular aggregate, thermodynamics, phase

behavior, pressure Fundamental Lecture⟩ “Physical Chemistry 2”(1.0), “Biophysical Chemistry 2”

(1.0)

Relational Lecture⟩ “Advanced Biophysical Chemistry”(0.5) Requirement⟩ Students are required to have a good understanding of undergraduate-

level physical chemistry and biophysical chemistry and related subjects. Notice⟩ [注意]

Goal⟩

1. To understand the thermodynamic treatments of molecular aggregates formed by bio-related substances.

2. To understand the nature for aggregates of bio-related substances and effects

of environmental variables on the aggregates.

Schedule⟩

1. Thermodynamics of molecular aggregates (1) thermodynamics of adsorption at interfaces 1: surface and interfacial tension

2. Thermodynamics of molecular aggregates (2) thermodynamics of adsorption at interfaces 2: treatments of adsorbed film

3. Thermodynamics of molecular aggregates (3) phase transitions of monolayers 1: insoluble monlayers 4. Thermodynamics of molecular aggregates (4) phase transitions of monolayers

2: soluble monolayers

5. Thermodynamics of molecular aggregates (5) thermodynamics of selfassociation 1: phase separation model

6. Thermodynamics of molecular aggregates (6) thermodynamics of selfassociation 2: mass action model

7. Thermodynamics of molecular aggregates (7) micelle formation in dilute solutions 1: critical micelle concentration

8. Thermodynamics of molecular aggregates (8) micelle formation in dilute solutions 2: affecting factors

9. Structures and functions of molecular aggregates (1) solution behavior 1: Krafft point

10. Structures and functions of molecular aggregates (2) solution behavior 2: cloud point

11. Structures and functions of molecular aggregates (3) phase behavior of bilayers 1: saturated phospholipids

12. Structures and functions of molecular aggregates (4) phase behavior of bilayers 2: unsaturated phospholipids

13. Structures and functions of molecular aggregates (5) pressure effect on bilayers 1: volume change of transition

14. Structures and functions of molecular aggregates (6) pressure effect on bilayers 2: pressure-induced bilayer interdigitation

15. Structures and functions of molecular aggregates (7) phase behavior of bilayer mixtures: domain formation

16. Summary, inquiry and report preparations Evaluation Criteria⟩ More than 80% percentage of attendance and reports (100%). Textbook⟩ To be distributed materials adequately in the class. Reference⟩ To be introduced in the class.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216731

Student⟩ Able to be taken by student of other department

Contact⟩

⇒ Matsuki (G607, +81-88-656-7513, matsuki@bio.tokushima-u.ac.jp) ���� (Office Hour: Friday 16:20-17:50)

Note⟩

⋄ When you take this class, it is necessary to do preparation for 2h and review for 2h everry 2h class for your understanding and taking credit. ⋄ Goal 1 is related to schedules 1-8 and goal 2 is related to schedules 9-15.